OpenStudy (anonymous):

help solving -2/5 / 2

OpenStudy (anonymous):

\[\frac{-1}{5}\]

OpenStudy (anonymous):

-0.2

OpenStudy (anonymous):

please explain steps

OpenStudy (anonymous):

-2 -2 --- ------- 5 = 2 * 5 --- 2

OpenStudy (anonymous):

thank you